# Densities Of Primes And Primitive Roots

### v12

@@ Abstract
Let $u\neq\pm 1,v^{2}$ be a fixed integer, let $p\geq 2$ be a prime, and let $\operatorname{ord}_{p}(u)=d\>|\>p-1$ be the order of $u\operatorname{mod}p$. This note provides an effective lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-1}$ for the number of primes $p\leq x$ with a fixed primitive root $u\text{ mod }p$ for all large numbers $x\geq 1$. The current results in the literature have the lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-2}$, and restrictions on the fixed primitive root to a subset of at least three or more integers. October 8, 2025
@@ 1 Introduction
A conditional proof of this result was achieved in [22], and simplified sketches of the proof appear in [29, p. 8], and similar references. The determination of the constant $\delta(u)\geq 0$ for a fixed integer $u\in\mathbb{Z}$ is an interesting technical problem, [22, p. 218], [26], [27], et alii. An introduction to its historical development, and its calculations is covered in [29, pp. 3–10], and [41].
@@
where $a_{0}=\prod_{p\geq 2}\left(1-p^{-1}(p-1)^{-1}\right)$ is Artin constant, and $B>1$ is an arbitrary number, was proved in [17] unconditionally, and refined in [41]. These works had shown that almost all admissible integers $u\in\mathbb{Z}-\left\{-1,1,v^{2}:v\in\mathbb{Z}\right\}$ are primitive roots for infinitely many primes. The number of exceptions is a subset of zero density in $\mathbb{Z}$. The individual quantity $\pi_{u}(x)$ in (1.2) can be slightly different from the average quantity in (1.3). The variations, discovered by the Lehmers using numerical experiments, depend on the primes decomposition of the fixed value $u$, see [22, p. 220], [29, p. 3], and similar references for the exact formula for the density $\delta(u)\geq 0$.
The Artin primitive root conjecture for functions fields was proved by Bilharz, see [39, Theorem 10.11], and the same conjecture for polynomials over finite fields was proved in [37]. However, there is no known infinite sequence of primes with a fixed primitive root. The current literature has results infinite sequences of primes with unknown primitive root in a small finite set. For example, in [18] it was proved that for a fixed primes triple $q>r>s\geq 2$, the subset of integers
@@
see [HB86]. In both of these results, the lower bound for the number of primes $p\leq x$ with a fixed primitive root in either of the subset $\mathcal{A}(q,r,s)$ or $\mathcal{B}(q,r,s)$ has the lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-2}$ for all large number $x\geq 1$.
@@ 2.1 Simple Characters Sums
where $v=\log u$ is the discrete logarithm of $u\neq 0$ with respect to some primitive root, and for some $k\geq 1$, see [28, p. 187], [33, p. 118], and [24, p. 271] for more details.
@@
for some $k\geq 1,\gcd(k,q)=1$, see [28, p. 187], [33, p. 118], and [24, p. 271]. The additive character sums are quite similar to Lemma 2.1.
@@ 3.2 Divisors Dependent Characteristic Function
The works in [8], and [45] attribute this formula to Vinogradov. The proof and other details on the characteristic function are given in [11, p. 863], [28, p. 258], [30, p. 18]. The characteristic function for multiple primitive roots is used in [6, p. 146] to study consecutive primitive roots. In [9], [25] it is used to study the gap between primitive roots with respect to the Hamming metric, et cetera. And in [45] it is used to prove the existence of primitive roots in certain small subsets $A\subset\mathbb{F}_{p}$. In [8] it is used to prove that some finite fields do not have primitive roots of the form $a\tau+b$, with $\tau$ primitive and $a,b\in\mathbb{F}_{p}$ constants. In addition, the Artin primitive root conjecture for polynomials over finite fields was proved in [37] using this formula.
@@ 3.3 Divisors Free Characteristic Function
It often difficult to derive any meaningful result using the usual divisors dependent characteristic function of primitive elements given in Lemma 3.2. This difficulty is due to the large number of terms that can be generated by the divisors, for example, $d\mid p-1$, involved in the calculations, see [11], [9] for typical applications and [29, p. 19] for a discussion.
@@ 5.1 Power Exponential Sum with Relatively Prime Index
A different approach to this result appears in [14, Theorem 6], and related results are given in [13], [19], [5], and [20, Theorem 1]. The upper bound given in Theorem 5.1 seems to be optimum. A different proof, which has a weaker upper bound, appears in [14, Theorem 6], and related results are given in [5], [13], [20], and [20, Theorem 1].
@@ 5.2 FFT of Power Exponential Sum with Relatively Prime Index
Let $p\geq 2$ be a large prime. If $\tau$ be a primitive root modulo $p$ and $0\neq a=o(p)$ is not a primitive root, then
@@
where $\delta>0$ is a small number and the implied constant is independent of $a$ and $b\in[1,p-1]$.
@@ 6 Evaluation Of The Main Term
([42, Lemma 1]) Let $x\geq 1$ be a large number, and let $\varphi(n)$ be the Euler totient function. Then
@@
A more general version of this Lemma is proved in [44], and related discussions are given in [29, p. 16]. The generalization to number fields appears in [23]. These results are ubiquitous in various results in Number Theory. Here, the logarithm integral is defined by
@@ 7 Estimate For The Error Term
Let $\psi(z)=e^{i2\pi z/p}$. By hypothesis $u\neq\tau^{n}$ for any $n\geq 1$ such that $\gcd(n,p-1)=1$. Rearrange the triple finite sum in the form
@@
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\left(\sum_{0<k\leq p-1}e^{-i2\pi\frac{uk}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\frac{k\tau^{n}}{p}}\right).$
Applying Lemma 5.1 to the inner sum and completing the upper bound of the finite sum return
@@ 8 Densities Of Primes And Primitive Roots
A formula for computing the density constant $\delta(u)\geq 0$, which is the density of primes with a fixed primitive root $u\neq\pm 1,v^{2}$, is derived in [22, p. 220]. This formula specifies the density as follows. Let $u=(st^{2})^{k}\neq\pm 1,v^{2}$ with $s$ squarefree, and $k\geq 1$, and let
@@
Information on the determination of the constant $\delta(u)\geq 0$, which is the density of primes with a fixed primitive root $u\neq\pm 1,v^{2}$, is provided in the next section and in [22, p. 220].
@@ 9 Application To Repeated Decimals
with $x_{i}\in\{0,1,2,\ldots,9\}$, was investigated by Gauss and earlier authors centuries ago, see [4] for a historical account, and [32] for recent developments. As discussed in [16, Articles 14-18], the period, denoted by $\operatorname{ord}_{p}(10)=d\geq 1$, is a divisor of p1p-1. The problem of computing the densities for the subsets of primes for which the repeating decimals have very large periods such as d=(p1)/2d=(p-1)/2, and d=p1d=p-1, is a recent problem. This note considers the following result.
@@
are also important in the calculation of the period modulo $p^{2}$, confer [RP96, p. 333], [10] for other details.

### v13

@@ Abstract
Let $u\neq\pm 1,v^{2}$ be a fixed integer, let $p\geq 2$ be a prime, and let $\operatorname{ord}_{p}(u)=d\>|\>p-1$ be the order of $u\operatorname{mod}p$. This note provides an effective lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-1}$ for the number of primes $p\leq x$ with a fixed primitive root $u\text{ mod }p$ for all large numbers $x\geq 1$. The current results in the literature have the lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-2}$, and restrictions on the fixed primitive root to a subset of at least three or more integers. November 15, 2025
@@ 1 Introduction
A conditional proof of this result was achieved in [25], and simplified sketches of the proof appear in [32, p. 8], and similar references. The determination of the constant $\delta(u)\geq 0$ for a fixed integer $u\in\mathbb{Z}$ is an interesting technical problem, [25, p. 218], [29], [30], et alii. An introduction to its historical development, and its calculations is covered in [32, pp. 3–10], and [44].
@@
where $a_{0}=\prod_{p\geq 2}\left(1-p^{-1}(p-1)^{-1}\right)$ is Artin constant, and $B>1$ is an arbitrary number, was proved in [20] unconditionally, and refined in [44]. These works had shown that almost all admissible integers $u\in\mathbb{Z}-\left\{-1,1,v^{2}:v\in\mathbb{Z}\right\}$ are primitive roots for infinitely many primes. The number of exceptions is a subset of zero density in $\mathbb{Z}$. The individual quantity $\pi_{u}(x)$ in (1.2) can be slightly different from the average quantity in (1.3). The variations, discovered by the Lehmers using numerical experiments, depend on the primes decomposition of the fixed value $u$, see [25, p. 220], [32, p. 3], and similar references for the exact formula for the density $\delta(u)\geq 0$.
The Artin primitive root conjecture for functions fields was proved by Bilharz, see [42, Theorem 10.11], and the same conjecture for polynomials over finite fields was proved in [40]. However, there is no known infinite sequence of primes with a fixed primitive root. The current literature has results infinite sequences of primes with unknown primitive root in a small finite set. For example, in [21] it was proved that for a fixed primes triple $q>r>s\geq 2$, the subset of integers
@@
see [24]. In both of these results, the lower bound for the number of primes $p\leq x$ with a fixed primitive root in either of the subset $\mathcal{A}(q,r,s)$ or $\mathcal{B}(q,r,s)$ has the lower bound $\{p\leq x:\operatorname{ord}_{p}(u)=p-1\}\gg x(\log x)^{-2}$ for all large number $x\geq 1$.
@@ 2.1 Simple Characters Sums
where $v=\log u$ is the discrete logarithm of $u\neq 0$ with respect to some primitive root, and for some $k\geq 1$, see [31, p. 187], [36, p. 118], and [27, p. 271] for more details.
@@
for some $k\geq 1,\gcd(k,q)=1$, see [31, p. 187], [36, p. 118], and [27, p. 271]. The additive character sums are quite similar to Lemma 2.1.
@@ 3.2 Divisors Dependent Characteristic Function
The works in [9], and [48] attribute this formula to Vinogradov. The proof and other details on the characteristic function are given in [14, p. 863], [31, p. 258], [33, p. 18]. The characteristic function for multiple primitive roots is used in [7, p. 146] to study consecutive primitive roots. In [10], [28] it is used to study the gap between primitive roots with respect to the Hamming metric, et cetera. And in [48] it is used to prove the existence of primitive roots in certain small subsets $A\subset\mathbb{F}_{p}$. In [9] it is used to prove that some finite fields do not have primitive roots of the form $a\tau+b$, with $\tau$ primitive and $a,b\in\mathbb{F}_{p}$ constants. In addition, the Artin primitive root conjecture for polynomials over finite fields was proved in [40] using this formula.
@@ 3.3 Divisors Free Characteristic Function
It often difficult to derive any meaningful result using the usual divisors dependent characteristic function of primitive elements given in Lemma 3.2. This difficulty is due to the large number of terms that can be generated by the divisors, for example, $d\mid p-1$, involved in the calculations, see [14], [10] for typical applications and [32, p. 19] for a discussion.
@@ 5.1 Power Exponential Sum with Relatively Prime Index
A different approach to this result appears in [17, Theorem 6], and related results are given in [16], [22], [6], and [23, Theorem 1]. The upper bound given in Theorem 5.1 seems to be optimum. A different proof, which has a weaker upper bound, appears in [17, Theorem 6], and related results are given in [6], [16], [23], and [23, Theorem 1].
@@ 6 Evaluation of the Main Term
([45, Lemma 1]) Let $x\geq 1$ be a large number, and let $\varphi(n)$ be the Euler totient function. Then
@@
A more general version of this Lemma is proved in [47], and related discussions are given in [32, p. 16]. The generalization to number fields appears in [26]. These results are ubiquitous in various results in Number Theory. Here, the logarithm integral is defined by
@@ 7 Estimate for the Error Term
Let $\psi(z)=e^{i2\pi z/p}$. By hypothesis $u\neq\tau^{n}$ for any $n\geq 1$ such that $\gcd(n,p-1)=1$. Rearrange the triple finite sum in the form of a discrete Fourier transform
@@
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\left(\sum_{0<k\leq p-1}e^{\frac{-i2\pi uk}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{\frac{i2\pi k\tau^{n}}{p}}\right).$
Applying Lemma 5.1 to the inner double sum and completing the upper bound of the finite sum return
@@ 8 Densities of Primes and Primitive Roots
A formula for computing the density constant $\delta(u)\geq 0$, which is the density of primes with a fixed primitive root $u\neq\pm 1,v^{2}$, is derived in [25, p. 220]. This formula specifies the density as follows. Let $u=(st^{2})^{k}\neq\pm 1,v^{2}$ with $s$ squarefree, and $k\geq 1$, and let
@@
Information on the determination of the constant $\delta(u)\geq 0$, which is the density of primes with a fixed primitive root $u\neq\pm 1,v^{2}$, is provided in the next section and in [25, p. 220].
@@ 9 Application To Repeated Decimals
with $x_{i}\in\{0,1,2,\ldots,9\}$, was investigated by Gauss and earlier authors centuries ago, see [5] for a historical account, and [35] for recent developments. As discussed in Articles 14-18 in [19], the period, denoted by $\operatorname{ord}_{p}(10)=d\geq 1$, is a divisor of p1p-1. The problem of computing the densities for the subsets of primes for which the repeating decimals have very large periods such as d=(p1)/2d=(p-1)/2, and d=p1d=p-1, is a recent problem. This note considers the following result.
@@
are also important in the calculation of the period modulo $p^{2}$, confer [41, p. 333], [DK11] for other details.
@@
10 Alternative Estimate for the Error Term
The altrrnative estimate for the error term $E(x)$ ocurring in the proof Theorem 1.1 is a recently new proof. It is significantly shorter than the original proof in Section 7.
Lemma 10.1.
$\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{\begin{subarray}{c}1\leq s\leq p-1\\ \gcd(s,p-1)=1\end{subarray}}\sum_{0<t\leq p-1}e^{\frac{i2\pi(\tau^{s}-u)t}{p}}=O\Bigg(xe^{-c_{4}\sqrt{\log x}}\Bigg),$ (10.1)
where $c>0$ is an absolute constant, for all sufficiently large numbers $x\geq 1$.
Proof.
Employing the characteristic function
[…]
of relatively prime integers to remove the double index in the first line of (10) and switching the order of summation in the second line yield
$E(x)$ $=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{\begin{subarray}{c}1\leq s\leq p-1\\ \gcd(s,p-1)=1\end{subarray}}\sum_{0<t\leq p-1}e^{\frac{i2\pi(\tau^{s}-u)t}{p}}$
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{1\leq s\leq p-1,}\sum_{0<t\leq p-1}e^{\frac{i2\pi(\tau^{s}-u)t}{p}}\sum_{\begin{subarray}{c}d\mid s\\ d\mid p-1\end{subarray}}\mu(d)$
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{d\leq p-1}\mu(d)\sum_{\begin{subarray}{c}1\leq s\leq p-1,\\ d\mid s\end{subarray}}\sum_{0<t\leq p-1}e^{\frac{i2\pi(\tau^{s}-u)t}{p}}$
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{d\leq p-1}\mu(d)\sum_{1\leq m\leq(p-1)/d,}\sum_{0<t\leq p-1}e^{\frac{i2\pi(\tau^{s}-u)t}{p}}.$
By hypothesis $\tau^{s}-u=\tau^{dm}-u\neq 0$ for any $s\geq 1$ such that $\gcd(s,p-1)=1$. Thus, evaluating the inner sum indexed by $t$ followed by the evaluation of the resulting inner sum indexed by $m$ yield
$E(x)$ $=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{d\leq p-1}\mu(d)\sum_{1\leq m\leq(p-1)/d}(-1)$
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{d\leq p-1}\mu(d)\left[\frac{p-1}{d}\right]$
$=$ $\sum_{x\leq p\leq 2x}\frac{p-1}{p}\sum_{d\leq p-1}\frac{\mu(d)}{d}-\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{d\leq p-1}\mu(d)\left\{\frac{p-1}{d}\right\}\,$
where the last line follows from the largest integer function $z]=z-\{z\}$ identity.
Applying Theorem 11.1 to the first double finite sum and Lemma 11.1 to the second double finite sum return
$E(x)$ $=$ $O\Bigg(\sum_{x\leq p\leq 2x}e^{-c_{1}\sqrt{\log p}}\Bigg)+\sum_{x\leq p\leq 2x}\frac{1}{p}\Bigg(-1+O\bigg(pe^{-c_{2}\sqrt{\log p}}\bigg)\Bigg)$ (10.7)
$=$ $O\Bigg(\sum_{x\leq p\leq 2x}e^{-c_{3}\sqrt{\log p}}\Bigg)+O\bigg(\frac{1}{\log x}\bigg)$
$=$ $O\Bigg(xe^{-c_{4}\sqrt{\log x}}\Bigg),$
where the last line follows from the prime number theorem on the short interval $[x,2x]$ and $c_{1},\ldots,c_{4}>0$ are constants. ∎
11 Average Orders of Mobius Functions
Theorem 11.1.
If $\mu:\mathbb{N}\longrightarrow\{-1,0,1\}$ is the Mobius function, then, for any large number $x\geq 1$, the following statements are true.
  1. (i)
    $\sum_{n<x}\mu(n)=O\left(xe^{-c\sqrt{\log x}}\right)$,
    unconditionally,
  2. (ii)
    $\sum_{n<x}\frac{\mu(n)}{n}=O\left(e^{-c\sqrt{\log x}}\right),$
    unconditionally,
where $c>0$ is an absolute constant.
Proof.
See [12, Equation 27.11.12], [11, p. 6], [36, p. 182], et alii. ∎
Lemma 11.1.
If $x\geq 1$ is a large number, $\mu$ is the Mobius function and $\{z\}$ is the fractional part function, then
$\sum_{n\leq x}\mu(n)\bigg\{\frac{x}{n}\bigg\}=-1+O\left(xe^{-c\sqrt{\log x}}\right),$ (11.1)
where $c>0$ is an absolute constant.
Proof.
Start with the identity
$1$ $=\sum_{n\leq x}\mu(n)\bigg[\frac{x}{n}\bigg]$ (11.2)
$=x\sum_{n\leq x}\frac{\mu(n)}{n}-\sum_{n\leq x}\mu(n)\bigg\{\frac{x}{n}\bigg\},$
see [2, Theorem 3.12]. Applying Theorem 11.1 to the first finite sum on the left side and rearranging it complete the verification. ∎